OpenStudy (anonymous):

Apples cost $2 per pound, and grapes cost $3 per pound. write and simplify an expression for the total cost if you buy 10 lbs of apples and grapes combined. (Use only the variable a)

OpenStudy (anonymous):

\[2a+3(10-a)\]

OpenStudy (anonymous):

apples=$2/pound grapes=$3/pound a pounds of apples g pounds of grapes Total is 10 pounds, so g=10-a cost=a*2+g*3=a*2+(10-a)*3=2a+10+3a=5a+10

OpenStudy (anonymous):

oops last line - should be: cost=a*2+g*3=a*2+(10-a)*3=2a+30-3a=30-a